In a factory, 30 noisy machines produce a decible rating of 90dB. How many machines must be removed aiming to decrease the decible level to 80 dB?
1
Expert's answer
2020-03-18T10:15:01-0400

If the intensity of sound produced by one machine is I, then the intensity of sound of 30 machines is 30I. In decibels, this intensity will be equal to:


"90=10\log_{10}{\frac{30I}{I_0}}"

"80=10\log_{10}{\frac{nI}{I_0}}"

"90-80=10=10\log_{10}{\frac{30I}{I_0}}-10\log_{10}{\frac{(30-x)I}{I_0}}"

"10=10\log_{10}{\frac{30}{30-x}}"

"x=27"

27 machines must be removed.